# Shaw Walters1∗  Sam Gao1,2∗  Shakker Nerd1  Feng Da1Warren Williams1 Ting-Chien Meng1Hunter Han2Frank He3Allen Zhang4Ming Wu5 Timothy Shen6Maxwell Hu7Jerry Yan8

### v1

@@ 1 Introduction
In this paper, we introduce Eliza, a pioneering open-source web3-friendly agentic operating system designed to bridge this gap. Eliza is the first of its kind, offering a platform that makes the deployment of web3 applications not only possible but also effortless. We emphasize that every aspect of Eliza is crafted as a regular Typescript program, ensuring that it remains under the full control of its users while also providing seamless integration with web3 functionalities. This includes, but is not limited to, reading and writing blockchain data, interacting with smart contracts, and much more.
@@ 3 Design Principles
Roughness is better  Given a fixed amount of engineering resources, and all else being equal, the time saved by keeping the internal implementation of Eliza simple can be used to implement additional features, adapt to new situations, and keep up with the fast pace of progress in the field of AI and web3. Therefore it is better to have a simple but slightly incomplete solution than a comprehensive but complex and hard to maintain design.
@@ 4.2 Frameworks
As an open source framework, Eliza should be compared with its rivals of the same kind: including RIG, G.A.M.E, ZerePy, Heurist, REI. As shown in Fig. 2, we collect the feedbacks from over 50+ AI researchers and senior blockchain developers to acquire their subjective assessment toward current trending web3 AI agent frameworks, it can be easily observed that Eliza outperforms other frameworks in terms of key indictors: model providers, chain compatibility, functionality and social media.
@@ 5 ElizaOS
As general frameworks often limited to its highly abstract low-level details and use cases, the direction moving from generic to specialize becomes evident as time goes on. Plus, it is well-known that AI and LLMs are fields that evolve rapidly, with new concepts and ideas emerging every week. Abstractions like LangChain or AutoGPT, which are built around a variety of emerging technologies, find it difficult to withstand the test of time with their framework design.
@@
Based on the philosophy derived from previous AI Agent frameworks, we build a highly controllable and well-orchestrated framework that primarily focus on the web3 industry, serves for a simple expression: sweeping away the hurdles for developers in turning the mighty AI Agents into life.
@@ 5.1.3 Providers
To draw an analogy, providers can be likened to the human perceptual system, with their primary function being to:

### v2

# Shaw Walters1∗  Sam Gao1,2∗  Shakker Nerd1  Feng Da1Warren Williams1 Ting-Chien Meng1Amie Chow1Hunter Han2Frank He3Allen Zhang4 Ming Wu5Timothy Shen6Maxwell Hu7Jerry Yan8

@@ 1 Introduction
In this paper, we introduce Eliza, a pioneering open-source web3-friendly agentic operating system designed to bridge this gap. Eliza is the first of its kind, offering a platform that makes the deployment of web3 applications not only possible but also effortless. We emphasize that every aspect of Eliza is crafted as a regular Typescript program, ensuring that it remains under the full control of its users while also providing seamless integration with web3 functionalities. This includes, but is not limited to, reading and writing blockchain data, interacting with smart contracts, and much more functionality.
@@ 3 Design Principles
Roughness is better  Given limited engineering resources and all else being equal, keeping Eliza’s internal implementation simple saves time for adding features, adapting to new situations, and keeping pace with advancements in AI and Web3. Therefore, it is better to have a simple but slightly incomplete solution than a comprehensive yet complex and hard-to-maintain design.
@@ 4.2 Frameworks
As an open source framework, Eliza should be compared with its rivals of the same kind: including RIG, G.A.M.E, ZerePy, Heurist, REI. As shown in Fig. 2, we collect the feedbacks from over 50+ AI researchers and senior blockchain developers to acquire their subjective assessment toward current trending web3 AI agent frameworks, it can be easily observed that Eliza outperforms other frameworks in terms of key indicators: model providers, chain compatibility, functionality and social media.
@@ 5 ElizaOS
As general frameworks often limited to its highly abstract low-level details and use cases, the direction moving from generic to specialization becomes evident as time goes on. Plus, it is well-known that AI and LLMs are fields that evolve rapidly, with new concepts and ideas emerging every week. Abstractions like LangChain or AutoGPT, which are built around a variety of emerging technologies, find it difficult to withstand the test of time with their framework design.
@@
Based on the philosophy derived from previous AI Agent frameworks, we have built a highly controllable and well-orchestrated framework that primarily focuses on the Web3 industry. This framework simplifies the process of bringing powerful AI agents to life by removing hurdles for developers.
@@ 5.1.3 Providers
To draw an analogy, providers can be likened to the human perceptual system. Their primary function is to: